KjSaif (kjsaif):

The total cost, c, in dollars, of b bottles of water is c = 1.1b. What is the constant of proportionality?

KjSaif (kjsaif):

In decimal form

OpenStudy (will.h):

Alright first we will need to find at least 3 values of each variable C values 1, 2, 3 Now find b values with respect to the values of C that wr made means that we will substitute in the expression the values of C that we created and then solve for b When C = 1 b will equal to 0.9090 When C = 2 b will equal to 1.8181 When C = 3 b will equal to 2.7272 So now we have C: 1, 2, 3 B: 0.9090, 1.8181, 2.7272 Set the corresponding values as C/B 1/0.9090= 1.1 2/1.8181 = 1.1 3/2.7272 = 1.1 That's the constant of proportionality

OpenStudy (will.h):

BTW slope is always constan of proportionality.

KjSaif (kjsaif):

thanks!

OpenStudy (will.h):

You welcome